INAF queue observing, 2015 Feb 12-13

Observer: F. Cusano, R. Carini
Telescope Operator: S. Allanson
Support Astronomer: Olga Kuhn on call
Instrument: MODS-LBC

Summary

At the sunset there are some cirrus and clouds.

Data Summary

Project_RUN Objects Filter ExpTime Notes
Israel A1 cats021950 dual grating 2 h COMPLETED
standard Hiltner600 dual + red grating    
Saracco B RDCS red grating 3 h seeing very variable COMPLETED
Mignoli MOS 20 J1411+12 red grating 2 h 30 last 15 m bad seeing
Mignoli LBC MOS20_red J1411+12 filter z-SLOAN 30 m  
Mignoli LBC_MOS20_red J1411+12 filter i-SLOAN 30 m  
Mignoli LBC_MOS20_red J1411+12 filter r-SLOAN 30 m  
Mignoli LBC_MOS20_red J1411+12 filter z-SLOAN 24 m  
sky flat LBC   filter i-SLOAN    

Details

01:32 TO open the dome

02:00 slew to Israel A1
  • acqMODS catsj021950_ls.acq mods1r.20150213.0002-0003
  • seeing 1.4"
  • the target is not visible, we get another image with exposure time =120 s mods1r.20150213.0004. also in this case is not visible
  • we get an image with t_exp=160 s mods1r.20150213.0005-0006 now it is ok
  • slit + target mods1r.20150213.0007

02:32 execMODS catsj021950_ls.obs mods1r.0008-0010, mods1b.0002-0004
  • seeing 1.2", airmass 1.22
  • 02:58 seeing 1.0"
  • during the exposition of the second image some cirrus have been passed.
  • last exposure with poor seeing and half at airmass higher than 1.5 (at the end of the exposure is 1.59)

04:00 we slew to the standard acqMODS Hiltner600.acq mods1r.0011-0012
  • airmass 1.17
  • execMods Hiltner600.obs dual grating mods1r.0013-0015, mods1b.0005-0007
  • red grating mods1r.0016-0018
  • seeing 1.8"

05:15 slew to Saracco B program
  • acqMODS mods.7.RDCSm1.acq
  • seeing about 1.4"
  • 05:29 we stop the acquisition because the seeing is 3"
  • 05:30 seeing 1.3", we continue the acquisition mods1r.0019-0022
  • 05:33 seeing 1"

05:37 execMODS mods.7.RDCSm1_8.obs
  • 05:40 seeing 4", we stop

06:08 we try acquisition again, seeing between 1.5" and 3" * acqMODS mods.7.RDCSm1.acq mods1r.0023-0026

06:24 execMODS mods.7.RDCSm1_8.obs mods1r.0027-0030
  • seeing variable between 1.1" and 1.5"
  • during the acquisition the seeing has reached values of 2.3" but also of 0.9" *07:31 seeing is decreasin, now varies between 1.1" and 1.5"

07:36 execMODS mods.7.RDCSm1_9.obs mods1r.0031-0034
  • seeing 1.2"
  • 08:02 seeing stable about 1.2"

08:47 execMODS mods.7.RDCSm1_9.obs mods1r.0035-0038
  • 09:12 dimm 2.4"
  • 09:33 dimm 12"

09:58 slew to Mignoli program MOS20
  • acqMODS mods.20..acq mods1r.0039-0041
  • seeing 1.2"

10:09 dofpia, /first, /X2, /redonly

10:14 execMODS mods.20.1.obs mods1r.0042-45
  • run 1_MOS20_z.ob. lbcr.20150213.103113.fits is not good, done during the offset of mods.
  • seeing 0.9"
  • 10:56 get one more image in the z filter.

11:01 dofpia, /X2, /redonly
  • seeing 0.8"

11:04 run 2_MOS20_i.ob
  • seeing 1"
  • 11.13 seeing 0.9"

11:26 execMODS mods.20.2.obs mods1r.0046-49
  • seeing 0.9"

11:52 dofpia, /X2, /redonly

11:58 run 3_MOS20_r.ob * seeing 1.0"

12.38 dofpia, /X2, /redonly

12:40 execMODS mods.20.3.obs mods1r.0050-51
  • seeing 1.1"

12:4 run 4_MOS20_z.ob
  • seeing 1.5"

*13.13 last frame in twilight

13.20 sky flat in i-SLOAN filter. 3 files are with I-Bessel for mistake

-- %USERSIG{RobertaCarini - 2015-02-13}%

Comments

 
Topic revision: r4 - 19 May 2015, FeliceCusano
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback